MICROWAVE SPECTRA AND STRUCTURES OF H_{2}O⋯AgF

\underline{S. L. STEPHENS}, N. R. WALKER, D. P. TEW AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K.

A Balle-Flygare FT-MW spectrometer coupled to a laser ablation source has been used to measure the pure rotational spectra of H_{2}O⋯AgF. Generation is via laser ablation (532 nm) of a silver rod in the presence of SF_{6}, argon, a low partial pressure of H_{2}O and the molecules are stabilized by supersonic expansion. The spectra of eight isotopologues have been measured. Rotational constants, B_{0} and C_{0}, and the centrifugal distortion constant, Δ_{J} have been determined. Isotopic substitutions are available at the silver, oxygen and hydrogen atoms. The spectra are consistent with a linear arrangement of oxygen, silver and fluorine atoms and the structure is either C_{2v} planar at equilibrium or C_{S} pyramidal but with a low potential-energy barrier to planarity such that the v = 0 and 1 states associated with the motion that inverts the configuration at the O atom are well separated.